OpenStudy (anonymous):

Algebra 2 question??

OpenStudy (anonymous):

Compare and contrast these four functions using complete sentences. Focus on the slope and y-intercept first, and then on any additional properties of each function.

OpenStudy (dumbcow):

first get all four into "y=mx+b" form so you can compare slopes and y-int

OpenStudy (anonymous):

How do i do that..

OpenStudy (dumbcow):

well for option 1: y-int is the initial value slope is rate of change (how much increase or decrease for each dog)

OpenStudy (dumbcow):

option 2: calculate slope using \[m = \frac{y_2 -y_1}{x_2 -x_1}\] then solve for y_int (b) \[y_1 = mx_1 +b\]

OpenStudy (dumbcow):

option 3: look at y_int on graph (0,5) calculate slope same as in option 2 using 2 points on graph (0,5), (2.5,0)

OpenStudy (anonymous):

okay, so can u help walk me threw each one and how to put it n sentences

OpenStudy (dumbcow):

what no , i just did walk you through each one

OpenStudy (anonymous):

but thers 4.. yuo only gave me 3

OpenStudy (dumbcow):

option 4 is given is "y=mx+b" form already

OpenStudy (anonymous):

its f(x) g(x) hx and j(x)

OpenStudy (anonymous):

ah okay,